OpenStudy (anonymous):

tan^-1(1/3) + tan^-1(1/5) +tan^-1 (1/7) = tan^-1 x.find x

OpenStudy (anonymous):

(tan^-1) 1/3 +1/5 8 4 -------- = tan^-1 -- = tan^-1 -- 1 - 1/15 14 7 tan^-1 4/7 + tna^-1 1/7 = tan-1 4/7 + 1/7 = tan^-1 35/45 =tan-1 7/9 --------- 1 - 4/49 hence x =7/9

OpenStudy (anonymous):

thank u....how did 1 -1/15 come in 1st line??

OpenStudy (anonymous):

tan^-1 a + tan^-1 b = tan^-1 a + b ---- 1 - ab This is the only formula I've used .. :)

OpenStudy (anonymous):

okie...got it...tnx a lot.. :)

OpenStudy (anonymous):

can u pls help me ?? tan(1/2 cos^-1 √5/3) = ? pls tell me which formula to use..

OpenStudy (anonymous):

now all u have to do is to convert cos into tan how do u do tht? now cos = b/h and tan =p/b use pytha. theorem and get the value for p :)

OpenStudy (anonymous):

p= 2?

OpenStudy (anonymous):

yeah u got tht right :)

OpenStudy (anonymous):

so tan(1/2 tan 2/√5) ?

OpenStudy (anonymous):

sry....tan(1/2 tan^-1 2/√5)